OpenStudy (trisarahtops):

PLEASE HELP Determine the slope of the graph of x^2 = ln(xy) at the point (1, e). 1 2 e e^3

zepdrix (zepdrix):

Hey :) So you'll need to differentiate implicitly, which means we'll forgo looking for an explicit function y(x) before differentiating. We'll just differentiate it as is. It will involve some product rule and stuff like that.

zepdrix (zepdrix):

\[\large\rm x^2=\ln(xy)\]Power rule on the left side,\[\large\rm 2x=\frac{d}{dx}\ln(xy)\]Do you remember your derivative for ln(x)?

OpenStudy (trisarahtops):

is it 1/x?

zepdrix (zepdrix):

Yes,\[\large\rm \frac{d}{dx}\ln(x)=\frac{1}{x}\]When we have more than just x inside of the log, we can generalize this rule with the `chain rule`:\[\large\rm \frac{d}{dx}\ln(stuff)=\frac{1}{stuff}(stuff)'\]

zepdrix (zepdrix):

\[\large\rm 2x=\frac{d}{dx}\ln(xy)\]Let's try that with our problem. Log(xy) becomes 1/(xy) when we differentiate it, ya? And then chain rule tells us we have to multiply by the derivative of the inner function (xy),\[\large\rm 2x=\frac{1}{xy}\cdot(xy)'\]

zepdrix (zepdrix):

Now we're taking the derivative of (xy). We have a product, so we need to apply our product rule.\[\large\rm (xy)'=x'y+xy'\]

zepdrix (zepdrix):

What do you think Sare? :O Too confusing?

OpenStudy (trisarahtops):

Yes, i think understand so far.

zepdrix (zepdrix):

When I write x', I mean \(\rm \dfrac{d}{dx}x\) So what is that derivative? :)

OpenStudy (trisarahtops):

1?

zepdrix (zepdrix):

Good.\[\large\rm (xy)'=(1)y+xy'\]\[\large\rm (xy)'=y+xy'\]our y' can not be simplied in a similar way. y'=dy/dx, the derivative function. This is what we're trying to solve for.

zepdrix (zepdrix):

\[\large\rm 2x=\frac{1}{xy}\cdot(xy)'\]So this is what we have so far,\[\large\rm 2x=\frac{1}{xy}\cdot(y+xy')\]Oh actually they gave us a point to plug in, I suppose we don't need to isolate our y' just yet. Let's plug in the coordinate they gave us first. It should make some of the math a bit easier on us. \(\large\rm (x,y)=(1,e)\)

zepdrix (zepdrix):

What do you get when you plug those x's and y's in? Wanna give it a shot? :)

OpenStudy (trisarahtops):

1/(1)(e) * (e+(1)(e)

zepdrix (zepdrix):

Close! :) We're not actually plugging anything for y' \(\rm \dfrac{1}{1\cdot e}\cdot(e+1y')\) And don't forget about the other side as well,\[\large\rm 2(1)=\dfrac{1}{1\cdot e}\cdot(e+1y')\]

zepdrix (zepdrix):

We've evaluated y' at the point (1,e). So this y' now represents the `slope` of the function at that point! Solving for y' will answer the question for us.

OpenStudy (trisarahtops):

wait i'm sorry why do you keep writing 2x. Doesn't the problem say x^2?

zepdrix (zepdrix):

That username is so cool -_- lol

zepdrix (zepdrix):

So what we did as a first step was: Take derivative of `both sides`.

OpenStudy (trisarahtops):

ohh okay that why it's 2x.

zepdrix (zepdrix):

\[\large\rm 2(1)=\dfrac{1}{1\cdot e}\cdot(e+1y')\]So your last step is to isolate the y'. Think you can do that? :)

OpenStudy (trisarahtops):

sure gimme one sec

OpenStudy (trisarahtops):

y=1?

zepdrix (zepdrix):

y' = 1 ? Hmm that doesn't look quite right :o

zepdrix (zepdrix):

Ignoring all of the 1's floating around, we currently have this,\[\large\rm 2=\dfrac{1}{e}(e+y')\]So what's a good first step in simplifying? Multiply both sides by e I suppose, ya?

zepdrix (zepdrix):

Which leads to,\[\large\rm 2e=e+y'\]

OpenStudy (trisarahtops):

now we divide right?

zepdrix (zepdrix):

No silly >.< We have e being `added` to y'. Inverse of addition? :)

OpenStudy (trisarahtops):

so subtract?

zepdrix (zepdrix):

Yes.

OpenStudy (trisarahtops):

e=2e-y or e=2-y

zepdrix (zepdrix):

We would like to isolate the y', we're trying to get everything on the left side of the equation. So we don't actually want to mess with the 2e, it's in the correct place. We'll undo the +e by subtracting e from each side.\[\large\rm 2e=e+y'\]\[\large\rm 2e-e=e-e+y'\]\[\large\rm 2e-e=y'\]Simply the left side.

OpenStudy (trisarahtops):

so y=e

OpenStudy (trisarahtops):

no wait 2?